a square poster has sides measuring 2 feet less than the sides of a square sign. If the difference between their areas is 24 square feet, find the lengths of the sides of the poster and sign.
Answer by checkley79(3341)

x^2=(x-2)^2+24
x^2=x^2-4x+4+24 cancel the x^2 terms.
-4x+28=0
-4x=-28
x=-28/-4
x=7 ans. for the sides of the sign.
7-2=5 ans. for sides the poster.
Proof:
7*2=5^2+24
49=25+24
49=49